shift summary, day 02/04/00


Subject: shift summary, day 02/04/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:53:13 CST


Shift Summary for day shift, 02/04/00

On Shift: Pat T., Leo B.

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 6.5
Hours of KAMI data: 6.5

Accounting of lost time: Access into NM2 to clear RF monitor from beam
                         after changing tageting angle to 12 mrad.

===============================================================
Runs ended this shift:
15758 KAMI_MUON
15759 KAMI_HADRON

===============================================================
Tapes brought to FCC this shift: none

Accesses during shift: 1) move BAG to west beam hole
                       2) lowered RF monitor in NM2

===============================================================
Description of Shift: KAMI muon run followed by targeting scans at 12 mrad

run spill time comment
----------------------------------------------------
15758 07:00 Day Shift starts.
             10:00 Rick F. shows up to change targeting angle from
                       15 mrad to 12 mrad.
             10:40 End run.

15759 10:50 Start run.
             11:00 It is clear that the beam is scraping something in
                       NM2. Accesses are made to move BAG to west beam hole
                       and to move RF monitor away from beam in NM2.
             11:20 End run and change tapes.

15761 13:20 Ready to start new run, but no beam yet.
             13:30 Beam returns, but does not seem to be hitting target,
                       MCR shuts off beam until Rick F. returns from KAMI
                       meeting. Problem was magnets which were not ramping.
                       Start target scans which continue through rest of
                       shift.



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:53:15 CST